OpenStudy (anonymous):

A stone is dropped from the top of a 450 m tower. (Acceleration due to gravity is -9.8 m/s^2. Ignore air resistance.) Find the distance of the stone above ground level at time t. A stone is dropped from the top of a 450 m tower. (Acceleration due to gravity is -9.8 m/s^2. Ignore air resistance.) Find the distance of the stone above ground level at time t. @Mathematics

OpenStudy (anonymous):

use h = 450 - 1/2(9.8)t^2 using h = ho +vot +1/2gt^2